OpenStudy (anonymous):

solve for z. x=1/2(22z-9w)

OpenStudy (anonymous):

hello u there?

OpenStudy (anonymous):

2x(22z-9w) = 1

OpenStudy (a_clan):

2x= (22z - 9w) 22z = 2x+ 9w z = (2x+9w)/22

OpenStudy (anonymous):

almost

OpenStudy (anonymous):

remember your dividing whats in the parenthesis

OpenStudy (anonymous):

so it will be 1 over that

OpenStudy (anonymous):

2x(22z-9w) = 1

OpenStudy (anonymous):

go from there

OpenStudy (anonymous):

i just cross multiplied to get that

OpenStudy (anonymous):

well, half of a cross multiply

OpenStudy (anonymous):

2x(22z-9w) = 1

OpenStudy (anonymous):

44xz - 18wx = 1

OpenStudy (anonymous):

44xz = 1+18wx

OpenStudy (anonymous):

z=(1+18wx)/44x

jhonyy9 (jhonyy9):

2x(22z-9w)=1 --- 44xz - 18xw =1 --- z=(1+18xw)/44x